# UM3: Unsupervised Map to Map Matching

### v1

@@ Abstract.
Map-to-map matching is a critical task for aligning spatial data across heterogeneous sources, yet it remains challenging due to the lack of ground truth correspondences, sparse node features, and scalability demands. In this paper, we propose an unsupervised graph-based framework that addresses these challenges through three key innovations. First, our method is an unsupervised learning approach that requires no training data, which is crucial for large-scale map data where obtaining labeled training samples is challenging. Second, we introduce pseudo coordinates that capture the relative spatial layout of nodes within each map, which enhances feature discriminability and enables scale-invariant learning. Third, we design an mechanism to adaptively balance feature and geometric similarity, as well as a geometric-consistent loss function, ensuring robustness to noisy or incomplete coordinate data. At the implementation level, to handle large-scale maps, we develop a tile-based post-processing pipeline with overlapping regions and majority voting, which enables parallel processing while preserving boundary coherence. Experiments on real-world datasets demonstrate that our method achieves state-of-the-art accuracy in matching tasks, surpassing existing methods by a large margin, particularly in high-noise and large-scale scenarios. Our framework provides a scalable and practical solution for map alignment, offering a robust and efficient alternative to traditional approaches. Code is open-sourced at https://github.com/LOGO-CUHKSZ/UM3.
Map-to-Map Matching, Unsupervised Learning, Graph Neural Networks
journalyear: 2025copyright: acmlicensedconference: The 33rd ACM International Conference on Advances in Geographic Information Systems; November 3–6, 2025; Minneapolis, MN, USAbooktitle: The 33rd ACM International Conference on Advances in Geographic Information Systems (SIGSPATIAL ’25), November 3–6, 2025, Minneapolis, MN, USAcopyright: noneccs: Information systems Information systems applications
@@ 2. Related Work
Graph matching aims to find correspondences between nodes and edges in two graphs, with applications in computer vision, bioinformatics, and network analysis. Early approaches, such as spectral matching (Leordeanu and Hebert, 2005), use graph Laplacians to identify correspondences but are sensitive to noise and structural differences. More recent methods incorporate deep learning to extract more dedicated features and optimize matching objectives (Zanfir and Sminchisescu, 2018; Yu et al., 2019; Sarlin et al., 2020). Other efforts have been devoted to designing advanced pipelines, such as combining with traditional non-differentiable combinatorial solvers (Rol´ınek et al., 2020), converting the graph matching problem into node classification problem on the association graphs (Wang et al., 2021b), and learning latent topology to enhance matching (Yu et al., 2021). While these techniques have shown promise, they often rely on supervised learning and struggle with scalability when applied to large graphs or heterogeneous data.
@@ 5.1. Experimental Setup
Our experiments are conducted on map data from three different regions: Boston (USA), Ichikawa (Japan), Shanghai (China), and Bremen (Germany), with statistics in Table 1. The dataset for Boston is obtained from a publicly available source (Bos, [n. d.]), the dataset for Ichikawa is acquired from an official government website (Okada et al., 2017), the dataset for Shanghai is collected from an industrial partner, and the dataset for Bremen is from OSM (Haklay and Weber, 2008). To evaluate the performance in real-world scenarios, we align these datasets with their corresponding regions in the latest OSM data. Specifically, in the Bremen dataset, we perform map-to-map matching between the 2014 and 2025 maps of a selected region, with both maps collected from OSM. Since the source and target maps originate from the same data source, their identifiers are consistent (e.g., unique road IDs in OSM), which makes it straightforward to obtain ground-truth correspondences. Due to differences in data sources, collection methods, and acquisition times, there exist inherent discrepancies between the collected road network data and the corresponding OSM data. These inherent differences make the dataset particularly challenging and realistic for evaluating map-to-map matching algorithms. The variety in geographic regions and data sources ensures the robustness of our method across different road network structures and data acquisition conditions.
@@ 5.2. Implementation Details
The other baseline methods were adapted from their official source code repositories to ensure consistency with their original implementations
@@ 6. Conclusion
Acknowledgements.
This work was supported by the National Key R&D Program of China under grant 2022YFA1003900 and a grant from MXNavi Co., Ltd

### v2

@@ Abstract.
Map-to-map matching is a critical task for aligning spatial data across heterogeneous sources, yet it remains challenging due to the lack of ground truth correspondences, sparse node features, and scalability demands. In this paper, we propose an unsupervised graph-based framework that addresses these challenges through three key innovations. First, our method is an unsupervised learning approach that requires no training data, which is crucial for large-scale map data where obtaining labeled training samples is challenging. Second, we introduce pseudo coordinates that capture the relative spatial layout of nodes within each map, which enhances feature discriminability and enables scale-invariant learning. Third, we design an mechanism to adaptively balance feature and geometric similarity, as well as a geometric-consistent loss function, ensuring robustness to noisy or incomplete coordinate data. At the implementation level, to handle large-scale maps, we develop a tile-based post-processing pipeline with overlapping regions and majority voting, which enables parallel processing while preserving boundary coherence. Experiments on real-world datasets demonstrate that our method achieves state-of-the-art accuracy in matching tasks, surpassing existing methods by a large margin, particularly in high-noise and large-scale scenarios. Our framework provides a scalable and practical solution for map alignment, offering a robust and efficient alternative to traditional approaches.
journalyear: 2025copyright: none
@@ 2. Related Work
Graph matching aims to find correspondences between nodes and edges in two graphs, with applications in computer vision, bioinformatics, and network analysis. Early approaches, such as spectral matching (Leordeanu and Hebert, 2005), use graph Laplacians to identify correspondences but are sensitive to noise and structural differences. More recent methods incorporate deep learning to extract more dedicated features and optimize matching objectives (Zanfir and Sminchisescu, 2018; Yu et al., 2019; Sarlin et al., 2020). Other efforts have been devoted to designing advanced pipelines, such as combining with traditional non-differentiable combinatorial solvers (Rolínek et al., 2020), converting the graph matching problem into node classification problem on the association graphs (Wang et al., 2021b), and learning latent topology to enhance matching (Yu et al., 2021). While these techniques have shown promise, they often rely on supervised learning and struggle with scalability when applied to large graphs or heterogeneous data.
@@ 5.1. Experimental Setup
Our experiments are conducted on map data from three different regions: Boston (USA), Ichikawa (Japan), Shanghai (China), and Bremen (Germany), with statistics in Table 1. The dataset for Boston is obtained from a publicly available source (1), the dataset for Ichikawa is acquired from an official government website (Okada et al., 2017), the dataset for Shanghai is collected from an industrial partner, and the dataset for Bremen is from OSM (Haklay and Weber, 2008). To evaluate the performance in real-world scenarios, we align these datasets with their corresponding regions in the latest OSM data. Specifically, in the Bremen dataset, we perform map-to-map matching between the 2014 and 2025 maps of a selected region, with both maps collected from OSM. Since the source and target maps originate from the same data source, their identifiers are consistent (e.g., unique road IDs in OSM), which makes it straightforward to obtain ground-truth correspondences. Due to differences in data sources, collection methods, and acquisition times, there exist inherent discrepancies between the collected road network data and the corresponding OSM data. These inherent differences make the dataset particularly challenging and realistic for evaluating map-to-map matching algorithms. The variety in geographic regions and data sources ensures the robustness of our method across different road network structures and data acquisition conditions.
@@ 5.2. Implementation Details
The other baseline methods were adapted from their official source code repositories to ensure consistency with their original implementations.